# Joint2Human: High-quality 3D Human Generation via Compact Spherical Embedding of 3D Joints

### 14 Dec 2023 v1

@@ 4.3 Multi-view Recarving Strategy
Since our training data is sampled along a fixed direction and the directionality of Fourier Occupancy Fields, there are a few artifacts along the direction orthogonal to the normal direction of the generated shape. To tackle this issue while maintaining muti-views consistency, we propose the multi-view recarving strategy to eliminate artifacts at edges. Different from the resampling in Chupa [21], we perform this process on 3D space and fuse them by calculating the blending occupancy field.
In detail, for each sampling process, we first sample a Gaussian noise $\tilde{z}_{t}\sim N(0,1)$. Then according to Eq. 4, we perform the complete denoising and decoding process to obtain the FOF $C_{init}$ and occupancy field $F_{init}$. Based on the $C_{init}$, we can get the initial human mesh $\mathcal{M}_{init}$, which can have some artifacts in other views. To tackle these artifacts, we transform the yaw of the body $\mathcal{M}_{init}$ orientation to get $\mathcal{M}_{\theta}$. After that, we turn the 3D mesh into FOF $C_{\theta}$ by sampling along the current orientation. With the $C_{\theta}$, we do the same thing as Sec. 4.1: leveraging the encoder $\mathcal{E}$ to get the latent code $\hat{z}$. After that, we re-execute the forward process of diffusion, adding a few steps of noise to $\hat{z}$. Then, we use the denoising model $\epsilon_{\theta}$ to denoise the noisy $\hat{z}$. As before, we pass the denoised latent code through the decoder $\mathcal{D}$ to get the FOF $C_{\theta}$. Meanwhile, we reconstruct the occupancy field $F_{\theta}$ from FOF $C_{\theta}$ by efficient Fourier inversion. At last, we blend the occupancy fields of different views and extract the 3D human mesh from the occupancy fields.
4.4 High-frequency Enhancer
For efficient human representation and modeling of data distribution, we only save low-frequency terms while using the Fourier occupancy field for representation. However, some abandoned high-frequency information is also important for human shape details. Therefore, based on the generation of low-frequency FOF feature in Sec. 4.1, we propose the High-Frequency enhancer to recover these precise details, as shown in Fig. 2. In detail, we learn a reference-based decoding network $\mathcal{D}_{h}$ based on the latent space and decoder from the auto-decoder claimed in Sec. 4.1 to produce high-frequency terms. To enhance the perception of latent-space contextual information, we introduce additional skip connections for each layer in $\mathcal{D}_{h}$, which is used to fuse the feature map from the relevant layers in the decoder $\mathcal{D}_{l}$. Then, we calculate the l2 loss between the predicted terms and high-frequency ground truth as the supervision to optimize $\mathcal{D}_{h}$. In this way, we achieve the map from the low-frequency domain to the high-frequency domain.
@@ 5.1 Experimental Setup
  • Datasets. We trained our model with THuman 2.0 [46], THuman 3.0 [41], 2k2k [14] and about 1500 high-quality meshes from commercial datasets. To ensure a fair comparison, we deviated from Chupa [21]’s settings and opted for a third-party dataset CustomHumans [15] as our test set. We do this because the sota methods’ training code is not publicly available and some of the datasets they use are not publicly available. For the training stage, we first sample 32 successive different angles of FOF along the yaw axis of the rotating human body to generate FOF feature maps for the same mesh. After that, we obtain the joint points for each mesh and calculate the compact spherical embedding of the 3D joint with the same sampling process of FOF.
  • Baselines. We compare our method with Chupa [21] and AG3D [9] as baselines. Chupa is the current state-of-the-art method for generating 3D human geometry. It can optimize the SMPL mesh by dual normal maps generated by the pose-guided diffusion model.
  • Metrics. We measure the quality of the generated human mesh by using the Fréchet Inception Distance (FID) [2] between the rendering normal maps [6, 21] and shading images [21, 40] of the generated meshes.
@@ 5.2 Implementation Details
We use an EMA rate of 0.9999 for all experiments. For the training stage of the FOF autoencoder, we trained it for 3 days on 8 NVIDIA A100 GPUs, with a batch size of 32. To represent and generate a high-quality 3D human geometry, the channels $N$ of the FOF feature maps need to be set to at least 32. However, it is very difficult to model high-dimensional and multi-channel data for diffusion models. To address this issue, following the latent diffusion model [38], we utilize an auto-encoder to compress the raw high-dimensional data space into a lower-dimensional latent space, which encodes each shape into a normal distribution. In detail, we adopted a VAE-like auto-encoder which contains the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. We present more details in the supplementary material.
@@
Figure 6: Our generated results guided by poses.
@@ 5.3 Generated Results
Fig. 5 shows various generated results by our method. It can be seen that our method can generate high-quality 3D humans with global structure, local details, and high resolution. Benefiting from our proposed compact spherical embedding of 3D joints, we can flexibly generate high-quality 3D humans guided by poses. Fig. 6 shows more generation results guided by poses. Although we have fixed the pose, our model still performs well in the diversity of generation. It can generate humans with different identities and costumes.
@@ 5.4 Comparison
Qualitative and Quantitative Results. We compare our method with the latest human generative models Chupa [21] and AG3D [9]. In detail, We conduct a quantitative comparison experiment with Chupa to assess the quality of the generated meshes. Apart from this, we conduct a qualitative comparison experiment with AG3D and Chupa. To ensure a fair comparison, we try to control that the input poses of Chupa and our method are the same as much as possible. However, AG3D is not friendly to fixed poses as input, so we used random poses in the test stage. The quantitative results for human generation are shown in Tab. 2. Our method outperforms the current SOTA methods and achieves lower FID on both images. Fig. 7 shows qualitative comparison results, comparing the results generated by AG3D, our results have more details. Our results not only look more natural but also have better view consistency and superiority for modeling loose clothing when compared to the generated result of other methods. So our method can ensure the global structure and local details simultaneously.
User Study. To better evaluate our model, we conduct a perceptual study to ask the users about their preferences for the generated body geometry, choosing among our approach and two other current state-of-the-art methods, Chupa[21] and AG3D[9]. The results of the survey are shown in Tab. 3. In the user study, we collected 123 answers, including 54 females and 69 males with different ages (6 users under 18, 113 users between 18 and 40, 2 users between 40 and 60, and 2 users above 60). .
In the study, we show the human generation results of AG3D (Method A), Chupa (Method B), and our method (Method C) in video form. Our study is divided into two sections, with a total of 7 cases. In the first stage, cases 1 through 5 are responsible for studying individual generation performance (each case has three questions that are the same from case to case, 15 questions in total). The three questions are about generating a 3D human shape with 1. Global structure; 2. Local detail; and 3. Overall impression. Each user who participated in the survey was asked to rank the results shown in the video from best to worst according to these three metrics. Tab. 3 demonstrates that our approach is the most popular throughout the user study.
In the second stage, we focus mainly on the generative diversity of the different approaches. Since AG3D cannot support guidance by the human body pose, we only study the performance of methods B, and C in this section. Different from the first stage, we show the user multiple results in the form of images with two cases, both asking them which method is better for diversity. Each case includes 6 results for both methods. In this section, a total of 70.73% of the users consider the diversity of our methods to be superior to Chupa (Method B).
@@ 5.5 Ablation Study
We have conducted extensive ablation studies to validate the components and settings of our pipeline. The quantitative comparisons use the same SMPL/pose guidance for evaluation. We also show and analyze different qualitative results to demonstrate the effectiveness of our module.
Analysis on Different Pose-guided Strategy. To validate the effectiveness of our pose-guidance, we train our model under different conditional information for pose-guidance. There are three types of conditional information in our study, SMPL [26], joint (our compact spherical embedding of 3D joints), and IUV [13]. We calculate the Fid scores under the combination of different pose conditions. The results are shown in Tab. 4. According to the results, it can be inferred that the compact spherical embedding of 3D joints can generally improve the generation quality.
@@
Ablation on Multi-view Recarving Strategy and High-Frequency Enhancer. The visual comparisons presented in Fig. 4 demonstrate that the recarving strategy significantly enhances the local details of the generated human. To further measure the ability of other modules to capture local details, we carry out comparative experiments under different modules and calculate the FID scores. The results are shown in Tab. 5.
@@ 6 Conclusion and Discussion
Conclusion. In this paper, we introduce Joint2Human, a novel and efficient method for directly generating detailed 3D human geometry using 2D diffusion models. We propose a compact spherical embedding of 3D joints, for flexible control and utilizable human prior. We also design a high-frequency enhancer and a multi-view recarving strategy to seamlessly integrate the details from different views into a uniform global shape, and hence guarantee both global structure and local details. Besides, our method can also generate high-quality 3D humans guided by text. Experimental results demonstrate that our method outperforms the state-of-the-art methods, making it ideal for advanced 3D applications.
Limitations. Although our method can produce results with various poses, it’s still a huge challenge for it to support extreme poses, such as stooping down or standing on the head. More failure cases are shown in the supplementary material.

### 06 Apr 2024 v2

@@ 3.4 Multi-view Recarving Strategy
Since our training data is sampled along a fixed direction and the directionality of FOF, there are a few artifacts along the direction orthogonal to the normal direction of the generated shape. We propose the multi-view recarving strategy to improve the geometry quality while maintaining muti-view consistency.
Different from the resampling in Chupa [22], we perform this process on 3D space and fuse them by blending the occupancy fields. In detail, we first perform the inference process of diffusion and high-frequency enhancer to obtain the FOF $C_{init}$ and the occupancy field $F_{init}$. Based on the $F_{init}$, we can get the initial human mesh $\mathcal{M}_{init}$, which may have some artifacts in other views. To tackle these artifacts, we rotate the $\mathcal{M}_{init}$ $\theta$ degrees along the yaw axis to get $\mathcal{M}_{\theta}$. After that, we convert the $\mathcal{M}_{\theta}$ into FOF $C_{\theta}$ along the current orientation. With the $C_{\theta}$, we do the same thing as described in Sec. 3.1: leveraging the encoder $\mathcal{E}$ to get the latent code $\hat{z}$. Then, we re-execute the forward process of diffusion, adding $T^{\prime}$ steps of noise to $\hat{z}$ and denoising the noise $\hat{z}$ with $\epsilon_{\theta}$. We pass the denoised latent code through the decoder to get the FOF $C_{\theta}$. Meanwhile, we can reconstruct the occupancy field $F_{\theta}$ from FOF $C_{\theta}$ by efficient Fourier inversion. At last, we blend the occupancy fields $\left\{F_{init},F_{\theta_{1}},F_{\theta_{2}}...\right\}$ of different views and extract the 3D human mesh from the occupancy fields. In our setting, we perform this procedure only once with the $\theta=\frac{\pi}{2}$. In this way, two orthogonal views are blended in a weighted-average approach. We settle on this setup under the trade-off between computational efficiency and geometric quality.
@@ 4.1 Experimental Setup
  • Datasets. We train our model with THuman 2.0 [49], THuman 3.0 [43], 2k2k [15] and about 1500 high-quality meshes from commercial datasets. To ensure a fair comparison, we deviated from Chupa [22]’s settings and used a third-party dataset CustomHumans [16] as our test set. This dataset wasn’t used to train any model. We do this because the Sota methods’ training code is not publicly available, and some of the datasets they use are not publicly available. For the training stage, we first sample 32 successive different angles of FOF along the yaw axis of the rotating human body to generate FOF feature maps for the same mesh. After that, we obtain the joint points for each mesh and calculate the compact spherical embedding of the 3D joint using the same FOF sampling process.
  • Baselines. We compare our method with Chupa [22] and AG3D [9] as baselines. Chupa is the current state-of-the-art method for generating 3D human geometry. We don’t make a quantitative comparison with AG3D because it uses 2D datasets, and the training code of AG3D is unavailable.
  • Metrics. We measure the quality of the generated human mesh by using the Fréchet Inception Distance (FID) [2] between the rendering normal maps [6, 22] and shading images [22, 42] of the generated meshes. We follow the settings of Chupa [22] and render the meshes into 18 views with $20^{\circ}$ yaw interval for calculating FID.
@@ 4.2 Implementation Details
To represent and generate a high-quality 3D human geometry, the channels of the FOF feature maps need to be set to at least 32. For the training stage of the FOF auto-encoder, we train it for three days on 8 NVIDIA A100 GPUs, with a batch size of 32. For the diffusion model training stage, we train it for eight days on 8 NVIDIA A100 GPUs, with a batch size of 64. The total number of time steps is set as $T=1000,T^{\prime}=200$ for the diffusion model in our pipeline. We present more details in the supplementary material.
@@
Figure 5: Our method can generate diverse results given a specific pose.
@@
Figure 7: The generated results with different pose-guidance strategies.
Figure 8: We present the side-view normal maps of the generated human. (a) illustrates the results without the High-Frequency Enhancer, while (b) illustrates the results using the High-Frequency Enhancer.
@@ 4.3 Generated Results
Fig. 4 shows various generated results by our method. Our method can generate high-quality 3D humans with global structure, local details, and high resolution. Benefiting from our proposed compact spherical embedding of 3D joints, we can flexibly generate high-quality 3D humans guided by poses as shown in Fig. 5. Our model performs well in the diversity of generations with the fixed pose as guidance. It can generate humans with different identities and costumes.
@@ 4.4 Comparison
Qualitative and Quantitative Results. We compare our method with the latest human generative models Chupa [22] and AG3D [9]. In detail, We conduct a quantitative comparison with Chupa to assess the quality of the generated meshes. We also conduct a qualitative comparison with AG3D and Chupa. To ensure a fair comparison, we try to make the input poses of Chupa and our method the same as much as possible. However, AG3D is not friendly to fixed poses of the results, so we used random poses in the test stage. The quantitative results for human generation are shown in Tab. 2. Our method outperforms the current SOTA methods and achieves better FID on both images. Fig. 6 shows qualitative comparison results. Our results gain the natural and detailed visual effects. Besides, we have better view consistency when compared to the generated results of other methods. In addition, our model can complete the generation process in less than one minute. It is advanced in loose clothing modeling and diversity generation. The discussion on running time and the visualized generation results are shown in the supplementary material.
User Study. To better evaluate our method with other state-of-the-art methods, we conduct a perceptual study to ask the users about their preferences of the three methods on the following three aspects: (1) Global structure, (2) Local detail, and (3) Overall impression. In the user study, we collected 123 answers from 54 females and 69 males of different ages (6 users under 18, 113 users between 18 and 40, 2 users between 40 and 60, and 2 users above 60). The results are shown in Tab. 3, demonstrating that our approach is the most popular throughout the user study. A total of 70.73% of the users consider our methods’ diversity superior to Chupa (Method B). More details about the user study can be found in the supplementary material.
@@ 4.5 Ablation Study
We have conducted extensive ablation studies to validate the components and settings of our pipeline. We use the same pose guidance in quantitative comparisons. We also show and analyze different qualitative results to demonstrate the effectiveness of our modules.
Ablation on Different Pose-guidance Strategies. To validate the effectiveness of our pose-guidance, we train our model under different conditional information for pose-guidance. There are three types of conditional information in our study, SMPL [29], CSEJ (our compact spherical embedding of 3D joints), and IUV [14]. We calculate the Fid scores under the different pose conditions. As shown in Tab. 4, the CSEJ can generally improve the generation quality. While CSEJ alone is not as good as IUV on this metric, the advantage of our CSEJ is enabling precise pose control. The FID only calculates the distance between two distributions but does not measure the ability to control the pose. Fig. 7 shows generated results using different guidance under the fixed pose; our CSEJ alone achieves better pose-control results than IUV.
@@
Ablation on Multi-view Recarving Strategies and High-Frequency Enhancer. Fig. 8 demonstrates the better visual effect of using High-Frequency Enhancer. The visual comparisons presented in Fig. 9 demonstrate that the recarving strategy significantly enhances the local details of the generated human. To further measure the ability of other modules to capture local details, we conduct comparative experiments under different modules and calculate the FID scores. The results are shown in Tab. 5.
@@ 5 Conclusion and Discussion
Conclusion. In this paper, we introduce Joint2Human, a novel and efficient method for directly generating detailed 3D human geometry using 2D diffusion models. We propose a compact spherical embedding of 3D joints for flexible control and the utilization of human prior. We also design a high-frequency enhancer and a multi-view recarving strategy to seamlessly integrate the details from different views into a uniform global shape, guaranteeing global structure and local details. Besides, our method can also generate high-quality 3D humans guided by text. Experimental results demonstrate that our method outperforms the state-of-the-art methods, making it ideal for advanced 3D applications.
Limitations. Although our method can produce results with various poses, supporting extreme poses is still a huge challenge, such as stooping down or standing on the head. More failure cases are shown in the supplementary material.
@@
Acknowledgements. This work was supported in part by National Key R&\&&D Program of China (2023YFC3082100), National Natural Science Foundation of China (62122058 and 62171317), and Science Fund for Distinguished Young Scholars of Tianjin (No. 22JCJQJC00040).